Break-Even Sales

Anheuser-Busch InBev Companies, Inc., reported the following operating information for a recent year:

Sales $4,352,000
Cost of goods sold $1,088,000
Selling, general and administration 544,000
$1,632,000
Income from operations $ 2,720,000*
*Before special items

In addition, assume that Anheuser-Busch InBev sold 34,000 barrels of beer during the year. Assume that variable costs were 75% of the cost of goods sold and 50% of selling, general, and administration expenses. Assume that the remaining costs are fixed. For the following year, assume that Anheuser-Busch InBev expects pricing, variable costs per barrel, and fixed costs to remain constant, except that new distribution and general office facilities are expected to increase fixed costs by $16,300.

a. Compute the break-even number of barrels for the current year. Round to the nearest whole barrel.
barrels

b. Compute the anticipated break-even number of barrels for the following year. Round to the nearest whole barrel.
barrels
Answer #1

For current year :

Variable cost Fixed cost
cost of goods sold 1088000*75%=816000 or 816000/34000 =$ 24 per barrel 1088000-816000= 272000
selling ,General and administration expense 544000*50%= 272000 or 272000/ 34000 =$ 8 per barrel 544000-272000=272000
Total $ 32 per barrel 544000
selling price 4352000/34000 =$ 128 per barrel

a)Breakeven point (number of barrels)= Fixed cost /(price-variable cost)

                                     = 544000/ (128- 32)

                                     = 544000 / 96

                                    = 5667 barrels

b)

Revised fixed cost= 544000+ 16300 = $ 560300

Breakeven point (number of barrels)= Fixed cost /(price-variable cost)

                                     = 560300/ (128- 32)

                                     = 560300 / 96

                                     = 5836.46 barrels (rounded to 5836)